OpenStudy (sphott51):

Which is a quadratic equation with a vertex of (4,-3_ and passes through point (6,2)? The answer is y= 5/4 (x-4^2 -3 .. I know that h=4 ad k= -3 so idk how to get the 5/4, so i need to show the work

OpenStudy (anonymous):

\[f(x)=a(x-4)-3\] solve for \(a\) by solving \[f(6)=2=a(6-4)-3\]

OpenStudy (anonymous):

you get \[2=a(6-4)^2-3\]

OpenStudy (anonymous):

whoa typos on the first two

OpenStudy (anonymous):

\[f(x)=a(x-4)^2-3\\ f(6)=a(6-4)^2-3=2\]

OpenStudy (anonymous):

\[4a-3=2\] solve for \(a\)

OpenStudy (sphott51):

Solve for a? how i do that?

OpenStudy (sphott51):

How to you get 4a-3=2?

OpenStudy (sphott51):

@satellite73

OpenStudy (sphott51):

Yeah I got it but how did you get 4a-3=2?